what principle must be invested at 12% if the simple interest earned after 2 years is $960?
Answer by nyc_function(2741)

Use the formula: I = PRT
We are looking for P.
960 = P(0.12)(2)
960 = P(0.24)
960/0.24) = P
4000 = P
The principal is $4000.
=======================

P is the principal (the initial amount you borrow or deposit)
r is the annual rate of interest (percentage or %)
I = interest
t = time in years
nycfunction@yahoo.com
Guido